Consider an economy having following values of Consumption, Investment, Government Spending, and Taxes 

C = 300 + 0.7Yd     I = 200 – 200i    G = 600    T = 100    

Derive the aggregate demand equation. Also, calculate the output in an economy. 



1
Expert's answer
2021-06-02T12:34:11-0400

AD = C + I + G +(X-M)

AD=aggregate demand

C=consumption

I=investment

G=government spending

X=total exports

M=total imports


AD= 300+0.7yd+200-200i+600+100

AD= 0.7yd+200i+1200


EO= consumption+investment+Government spending + Taxes

EO= 300+200+600+100= 1200